OpenStudy (anonymous):

Three friends all have ages that are consecutive integers. The sum of their ages is 180. What is the age of the oldest friend?

OpenStudy (anonymous):

(>u<) it was 61

OpenStudy (ipwnbunnies):

^^ That's not correct. .-. Let 'x' be the youngest friend. x+1 is the second oldest. x+2 is the oldest. Consecutive ages. x + (x+1) + (x+2) = 180 Solve for x. Find the age of the oldest. c:

OpenStudy (ipwnbunnies):

I meant the 180/3 was not correct. .~.

OpenStudy (anonymous):

THANK YOU LOL. Can you take a look at this one maybe? :) Jane is going to fence in her back yard. She has purchased 100 feet of fencing and knows that she wants to fence in a rectangular area where one side will be the back of her house. She knows her house is 30 feet across the back. Which of the following is the equation that Jane can use to figure out how far back from the house she can fence in?

OpenStudy (anonymous):

A. 30x=100 B. x + x + 30 = 100 C. x+30=100 D. 30(x+30)=100

OpenStudy (ipwnbunnies):

Let's draw a picture. ^_^

OpenStudy (anonymous):

carry the one

OpenStudy (ipwnbunnies):